OpenStudy (anonymous):

find the vol of a sphere whose surface area is 16pi cm

OpenStudy (anonymous):

256/3=83.33

OpenStudy (anonymous):

my options are 30pi/3 38pi/3 32pi/3 and 12pi/3

OpenStudy (anonymous):

4/3(2)^3*pi=32 pi/3

OpenStudy (anonymous):

4(r)^2 pi=16 pi r=2 vol=4/3(2)^3 pi=32pi/3